### v1

# Keyword-driven Retrieval-Augmented Large Language Models for Cold-start User Recommendations

Abstract.
Recent advancements in Large Language Models (LLMs) have shown significant potential in enhancing recommender systems. However, addressing the cold-start recommendation problem, where users lack historical data, remains a considerable challenge. In this paper, we introduce KALM4Rec (Keyword-driven Retrieval-Augmented Large Language Models for Cold-start User Recommendations), a novel framework specifically designed to tackle this problem by requiring only a few input keywords from users in a practical scenario of cold-start user restaurant recommendations. KALM4Rec operates in two main stages: candidates retrieval and LLM-based candidates re-ranking. In the first stage, keyword-driven retrieval models are used to identify potential candidates, addressing LLMs’ limitations in processing extensive tokens and reducing the risk of generating misleading information. In the second stage, we employ LLMs with various prompting strategies, including zero-shot and few-shot techniques, to re-rank these candidates by integrating multiple examples directly into the LLM prompts. Our evaluation, using a Yelp restaurant dataset with user reviews from three English-speaking cities, shows that our proposed framework significantly improves recommendation quality. Specifically, the integration of in-context instructions with LLMs for re-ranking markedly enhances the performance of the cold-start user recommender system.
Restaurant Recommendation, Cold-start User Recommendation, Large Language Model, Prompt Tunning.
isbn: 978-1-4503-XXXX-X/18/06ccs: Information systems Retrieval models and rankingccs: Information systems Language modelsccs: Information systems Learning to rank
1. Introduction
Recommender systems play a pivotal role in assisting users in navigating the vast array of available choices in the digital world. However, a significant and ongoing challenge in this field is addressing the issue of cold-start users. These are users who are new to the platform and therefore have little to no interaction history. The lack of data makes it difficult for the system to generate accurate and personalized recommendations. Conventional Collaborative Filtering (CF) such as BPR (Rendle et al., 2009), NCF (He et al., 2017) and LightGCN (He et al., 2020) or generative methods (Liang et al., 2018; Wang et al., 2023b; Liu et al., 2023) struggle to suggest relevant items effectively for these new users due to the lack of detailed preference information. While user-user content-based algorithms offer a solution by relying on user features to find similar users and recommend positively interacted items, this approach raises privacy concerns. Overall, there remains a significant gap in research addressing the needs of cold-start users, highlighting the necessity for further exploration in this area.
Figure 1. A pictorial illustration of our Keyword-driven Retrieval-Augmented Large Language Models for cold-start users recommendation (KALM4REC).
Meanwhile, Large Language Models (LLMs) have garnered considerable interest for their potential to enhance recommender systems, showcasing significant proficiency in natural language understanding, generation, and complex reasoning. Leveraging their robust architecture and extensive training on vast datasets, several studies employ LLMs to comprehend and address user requirements throughout the recommendation process (Gao et al., 2023; Wang et al., 2023a; Li et al., 2024; Wu et al., 2023). Moreover, researchers have explored LLMs’ potential to alleviate the cold-start recommendation problem through in-context learning (He et al., 2023). With LLMs’ adaptability to new tasks, prompts are often tailored to incorporate user and item information, facilitating effective adaptation and performance enhancement (Sanner et al., 2023; Wang et al., 2024).
Challenge. Despite their excellent capacities, existing LLMs suffer from several limitations when applied to recommender systems, especially for cold-start scenarios. For one, lacking comprehensive knowledge in a specific domain may result in non-factual outputs. While fine-tuning smaller LLMs could potentially reduce such hallucinations, this solution is often impractical due to the significant resources required (Mialon et al., 2023). Furthermore, directly incorporating user/item information can be costly in terms of token usage and restricts input length due to limited context length.
Approach. For recommender systems on the restaurant domain, user reviews offer valuable insights into their opinions about various businesses. Modeling these personal reviews is beneficial for comprehending users’ preferences and capturing the characteristics of restaurants (Vu et al., 2020). For cold-start users who lack historical data, a practical strategy is to ask them to provide a few keywords describing their preferences. This approach helps create an initial user profile with minimal effort and without compromising their privacy. Furthermore, using keywords instead of entire reviews to prompt LLMs can significantly enhance the efficiency and accuracy of the recommendations while minimizing token usage and computational costs.
To this end, we present a novel framework called Keyword-Driven Retrieval-Augmented Large Language Models for Cold-Start User Recommendations (KALM4REC). Our framework is built upon two essential components centered around keywords: candidates retrieval focusing on retrieving relevant restaurants, and LLM-ranker which leverages LLM to re-rank the retrieved candidates (illustrated in Figure 1). User reviews reflect user preferences but often contain noise, potentially leading to inaccurate representations of user preferences. By exclusively leveraging extracted keywords, we retain the contextual essence of the reviews while still describing the user profile well. Utilizing keywords also allows us to optimize the prompts by reducing noise, thereby enhancing their effectiveness while still ensuring sufficient information of users and restaurants in LLMs, effectively addressing cold-start issues. Essentially, KALM4REC, grounded in sets of keywords, aims to narrow down the candidate set relevant to predefined cold-start preferences. The LLM utilizes its contextual understanding and reasoning capabilities to generate a ranked list of restaurants for recommendations. Our contributions can be summarized as follows:
  • To the best of our knowledge, KALM4REC stands as the pioneering framework to employ keyword as the focal point of both the candidate retrieval and LLM-based re-ranking modules in recommender system. By prioritizing keywords, our augmented LLMs framework effectively addresses the significant challenge of hallucination outcome and optimizing prompt design, resulting in promising performance for cold-start users recommendation.
  • We propose approaches that take advantages of the extracted keywords to obtain most possible restaurants for cold-start users. Furthermore, we designed prompting templates to enable LLMs work efficiently for re-ranking tasks.
  • Finally, we conduct comprehensive experiments to validate and demonstrate the effectiveness of KALM4REC, compared to previous baseline methods for the cold-start user recommendation challenge.
2. Related works
Cold-start Recommendations. A recommendation system that aims to provide suggestions to new users with little or no historical data is known as a cold-start recommendation system. These systems play a critical role in businesses and services that require recommendations for new users or products. To address the challenge of generating precise recommendations for such cases, various approaches have been proposed. While collaborative filtering (CF) (He et al., 2017; Wang et al., 2019a) has been effective in traditional recommendation systems, it struggles with the cold-start scenario that arises when dealing when new users or items onboard the system. Previous CF-based methods (Bianchi et al., 2017; Gao et al., 2018) rely on data augmentations to incorporate users’ and items’ information into CF to deal with this problem. DropoutNet (Volkovs et al., 2017) employs dropout to directly train the model for cold-start issues, avoiding the need to aggregate all content-based information into objective terms. (Zhao et al., 2022) utilizes latent variables to develop a probabilistic approach for learning a distribution over the side information of items to warm up cold item ID embedding. (Wang et al., 2019b) consider knowledge graph as the source of side information to assist recommendation task. Although these methods show improvements in cold-start problems, they rely heavily on auxiliary information which usually causes privacy concerns.
LLMs for Recommendations. Meanwhile, Large Language Models (LLMs) have attracted significant interest for their ability to improve recommender systems, demonstrating notable skill in natural language understanding, generation, and complex reasoning. Capitalizing on their robust architecture and extensive training on vast datasets, various studies have utilized LLMs to understand and meet user needs throughout the recommendation process (Gao et al., 2023; Wang et al., 2023a; Li et al., 2024; Wu et al., 2023). Additionally, researchers have investigated the potential of LLMs to address the cold-start recommendation problem through in-context learning (He et al., 2023). Given LLMs’ adaptability to new tasks, prompts are often customized to incorporate user and item information, thereby enhancing their effectiveness and performance (Sanner et al., 2023; Wang et al., 2024). Hou et al. (2024) delves into the exploration of LLMs’ capabilities as zero-shot rankers for sequential recommendation. They achieve this by formulating recommendation as a conditional ranking task based on sequential interaction histories, employing meticulously crafted prompt templates, and conducting experiments on some benchmark datasets. Their findings demonstrate the promising zero-shot ranking power of LLMs, surpassing the performance of traditional models. Sanner et al. (2023); Wang et al. (2024) have investigated the inclusion of user and item information in prompts to address cold start challenges. However, none of those previous work have explored the use of keywords to augment LLMs in recommending to cold-start users.
Figure 2. KALM4REC begins with the Keyword Processing module, where we extract noun phrases as keywords from user reviews to construct their profiles. For cold-start users, candidate restaurants are generated using a retrieval model based on their selected terms. These candidates are then passed to the Prompt Generation stage, where information about user preferences and restaurant characteristics is introduced. Finally, leveraging their pre-trained knowledge, LLMs capture user expectations and perform re-ranking accordingly.
3. Proposed Framework
In this section, we present the details of our keyword-driven framework, KALM4REC, designed to tackle the cold-start recommendation problem. KALM4REC unfolds in two primary stages: (1) Candidates retrieval, where meaningful noun phrases are extracted from collected reviews to form word sets representing users and restaurants, and subsequently utilized for queries from cold-start users to obtain potential candidates. This phase employs our retrieval models, leveraging a collaborative filtering mechanism with keywords at its core to rank potential candidates; and (2) LLM-based candidates re-ranking, where an LLM is employed as a ranker to re-rank the retrieved candidates.
3.1. Problem Formulation
Given the set of users $\mathcal{U}$ and the set of restaurants $\mathcal{R}$, the interaction between a user $u\in\mathcal{U}$ and a restaurant $r\in\mathcal{R}$ is presented as a set of reviews =$\mathcal{I}=\{i_{u,r}\}$, where each entry $i_{u,r}$ represents a review if user $u$ has a review for restaurant $r$. The task involves recommending relevant restaurants to a new user $u_{c}$ (cold-start user). The new user can opt to declare his/her preference via a set of pre-defined keywords $k_{u_{c}}$. Simultaneously, user reviews also contain “keywords” $\mathcal{K}=\{k_{u,r}\}$, where each entry $k_{u,r}$ is a set of keywords extracted from $i_{u,r}$. We denote $k_{u}$ as a set of keywords extracted from all user $u$’s reviews, and $k_{r}$ as a set of keywords extracted from all reviews written for restaurant $r$. The problem of keyword-driven cold-start user recommendation is defined as follows. Given a set of keywords $k_{u_{c}}$ from a cold-start (new) user, return a ranked list of relevant restaurants $R_{u_{c}}$ where $R_{u_{c}}\subset\mathcal{R}$. In the following, we now present KALM4REC with the details of the two stages.
3.2. Candidates Retrieval
In this stage, we present methodologies for retrieving potential candidates that include Content-based Recommendation by Keywords (CBR) and Message Passing on Graph (MPG). For CBR, we utilize deep neural networks to learn the interaction function from sets of keywords representing users and restaurants. To accommodate the arbitrary order of keywords within sets, we adopt a pooling layer based on an attention mechanism. For MPG, we construct a heterogeneous graph comprising nodes representing keywords and restaurants, with two types of edges defined, and introduce a scheme for estimating the scores of the connections between keywords and restaurants.
3.2.1. Content-based Recommendation by Keywords (CBR)
Motivated by the effectiveness of existing works (Chen et al., 2022), we generate the embedding for the user and restaurants from its keyword representation. For CBR, profile of all users includes cold-start user is presented as sequence of keywords.
Given the sequence of keywords for user $u$: $E_{u}=[E_{u}^{1},E_{u}^{2},...,E_{u}^{T}]\in\mathbb{R}^{T\times d}$, where $T$ denotes the number of keywords for the user and $d$ is the dimension size of the keyword vectorized from sBERT (Reimers and Gurevych, 2019). The recommendation problem can be abstracted as $\hat{y}_{ur}=f(u,r|\Theta)$, where $f$ is a function learned to estimate the score for the user-restaurant pair. Let $p_{u}$ and $q_{r}$ denote the latent vectors for user $u$ and restaurant $r$, learned by $\Theta$. The interaction can be estimated as the inner product of $p_{u}$ and $q_{r}$:
(1) $\hat{y}_{ur}=f(u,r|p_{u},q_{r})=p_{u}^{T}q_{r}$
where $p_{u}$ and $q_{r}$ are learned to encode the user profile and restaurant profile, respectively, through a set pooling mechanism based on an attention mechanism (Safari et al., 2020) denoted as $G(E_{u})$.
(2) $G(E_{u})=\sigma(W_{c}E_{u}^{T})E_{u}$
where $\sigma$ indicates the softmax layer, and $W_{c}\in\mathbb{R}^{d}$ inferring a trainable parameter. $G(E_{u})$ can be interpreted as a dot product attention, where the keys and the values correspond to the same representation, and the query is only a trainable parameter. Therefore, $G(E_{u})$ represents a weighted average of the encoder sequence of features. The learned features are expected to fully represent the user profile.
For optimization, we adopt Bayesian Personalized Ranking (BPR) (Rendle et al., 2009), where $r_{p}$ is the restaurant that user $u$ prefers over $r_{n}$. The model is optimized by:
(3) $\sum_{(u,r_{p},r_{n})\in\mathcal{D}_{S}}\ln\sigma(\hat{y}_{u,r_{p},r_{n}})-\lambda_{\Theta}\|\Theta\|^{2}$
where $\hat{y}_{u,r_{p},r_{n}}$ represents the relation between user $u$ and restaurants $r_{p}$ and $r_{n}$; $\sigma$ is the logistic sigmoid function; $\lambda_{\Theta}$ are model regularization parameters.
For inference, the process is similar to training, as both stages involve encoding sets of keywords.
3.2.2. Message Passing on Graph (MPG)
Utilizing CBR alone is not sufficient to provide satisfactory embeddings for Collaborative Filtering. The primary deficiency lies in the embedding function’s failure to encode crucial collaborative signals latent in keyword-restaurant and keyword-user interactions. Drawing inspiration from LightGCN (He et al., 2020), which leverages structured information, we propose a model based on a tripartite graph structure (user - keyword - restaurant). Specifically, we construct the graph using training data, with edges established for (user - keyword) and (keyword - restaurant) interactions when a user uses a keyword to review and a restaurant’s review contains that keyword. Information is passed through the edges in the graph, and node information could be generated as follows:
(4) $q_{r}=AGG(q_{w},w\in k_{r})$
where $AGG$ is a function that aggregates information from neighboring nodes, and similarly for $q_{w}$.
However, due to the challenge posed by a large number of nodes, reaching millions, we propose an unsupervised learning model. First, we adopt an aggregator $(AGG)$ based on a simple weighted sum (similar to the work from LightGCN (He et al., 2020)), but without trainable parameters to determine the weights.
Figure 3. Illustrate graph construction with connections between users, keywords, and restaurants. Green nodes represent keywords from the training data, while dashed lines indicate connections established via nearest neighbors for keywords not in the data. Edge brightness reflects the importance score.
For a restaurant, the message is generated by multiplying the information of the keyword (source node) by its edge weight, then summing up the messages to update as follows:
(5) $q_{r}=\sum_{r\in\mathcal{R},w\in k_{r}}e^{w}*a^{w}_{r}$
where $e^{w}$ denotes feature of keyword nodes $w$; $a^{w}_{r}$ denote the connection between node $w$ and node $r$. Specifically, the edge between a keyword node and a restaurant can be represented by a value indicating the importance of a keyword to a restaurant, denoted as $\mathcal{A}_{r}^{w}$. We introduce a scheme named TF-IRF to measure the importance score of a keyword to a restaurant, which is similar to the TF-IDF idea. The score assigned to the edge established by a restaurant $r$ and a keyword $w$ is denoted as $s^{w}{r}$ and defined as:
(6) $a_{r}^{w}=tf_{r}^{w}*irf^{w}$
in which:
  • $tf_{r}^{w}=\frac{f_{r}^{w}}{q_{w}}$, where $f_{r}^{w}$ represents the number of times the term $w$ appears in the set of keywords for restaurant $r$, and $q_{w}$ indicates the total number of times the term $w$ is used in the entire training data.
  • $irf^{w}=\log\left(\frac{|\mathcal{R}|}{f^{w}}\right)$, where $f^{w}$ denotes the number of restaurants containing the term $w$.
The scores of all restaurants $\mathcal{S}$ can be rewritten in matrix form as:
(7) $\mathcal{S}=\mathcal{M}_{w}\times\mathcal{A}_{r}^{w}$
where $\mathcal{S}\in\mathbb{R}^{1\times n},\mathcal{A}_{r}^{w}\in\mathbb{R}^{w\times r}$, and $\mathcal{M}_{w}\in\mathbb{R}^{1\times w}$ denotes the occurence matrix represent the connection between cold-start user and their selected keywords. Then, the top-k possible restaurants are provided by:
(8) $C_{k}=\operatorname*{argmax}_{k=20}(\mathcal{S})$
Notably, no training is needed for this approach, only graph construction. If cold-start users select keywords absent in the training data, we replace them with the most similar keywords in semantics from the training data. To achieve this, we employ a pretrained BERT model (Reimers and Gurevych, 2019) to vectorize all keywords, ensuring that semantically similar keywords are close to each other. The most similar keyword then searched by using nearest neighbor model [^1] https://scikit-learn.org/stable/modules/neighbors.html.
3.3. LLM-based Candidates Re-ranking
[…]
Input: Please suggest the 15 most suitable restaurants for me from the candidate set that I will visit them, according to the user and candidate restaurant keywords I provided above.
Output: Must include 15 restaurants in the candidate restaurant set, not keyword. Desired format is string: restaurant_id_1, restaurant_id_2, … We utilize LLMs to re-rank candidates for each user, followed by natural language instructions. Additionally, leveraging their reasoning and generation abilities, we incorporate information about user and restaurant candidates into the instructions to make LLMs aware of user preferences. Like the retrieval model, user and restaurant information is represented by keywords. We also include sentences in Template $\mathcal{T}$ to trigger the recommender abilities of LLMs (Hou et al., 2024) and to describe the task instructions to the model:
Following, we propose a general restaurant recommendation prompt framework using keywords, consisting of (1) user keywords; (2) candidate restaurant sets; and (3) restaurant keyword sets. This prompt framework is denoted as [$\mathcal{H}_{z}$] and [$\mathcal{H}_{f}$] represent for zero-shot and few-shot pattern, respectively.
3.3.1. Zero-shot $\mathcal{H}_{z}$
[…]
The candidate restaurant set for me is enclosed in square brackets, with the restaurants separated by commas (Format: [restaurant_id_1, restaurant_id_2,…]) is: {candidate restaurant set}
Keywords associated with candidate restaurants have the following form: restaurant_id_1 (keyword 1, keyword 2,…) are {restaurant keywords sets}.
Since large language models trained on vast, information-rich datasets, zero-shot prompts require meaningful keywords for LLMs to process information and deliver high-quality re-ranking results. When keywords associated with the user are given, LLMs can grasp and anticipate the user’s interests. Therefore, the significance of keywords yields valuable insights. In the same way, the LLMs also comprehends the characteristics of restaurants, enabling it to recommend that are relevant and appealing to users based solely on the keywords provided about the restaurant.
3.3.2. Few-shot [$\mathcal{H}_{f}$]
[…]
# Repeat $\textbf{i}\in\{1,..,k\}$:
Example $i$:
There are the keywords that user often mention when wanting to choose restaurants: {user keywords}.
The candidate restaurant set for user is enclosed in square brackets, with the restaurants separated by commas (Format: [restaurant_id_1, restaurant_id_2,…]) is: {candidate restaurant set}
Keywords associated with candidate restaurants have the following form: restaurant_id_1 (keyword 1, keyword 2,…) are {restaurant keywords sets}.
You should rank a list of recommendations for user as follows: {restaurant_id_1, restaurant_id_2, …}
Pattern: [ $\mathcal{H}_{z}$ ]
The pattern of few-shot prompts [$\mathcal{H}_{f}$] above comprises examples and the information in [$\mathcal{H}_{z}$] then filled into Template $\mathcal{T}$. The keyword candidates are ordered in our prompts to preserve the order from retrieval models, which is based on TF-IRF scores.
Examples are generated from users in the training set. For each example, candidates are selected based on the number of overlapping keywords to a chosen user. The ranked list for each user is obtained by sorting the rating of the candidates.
4. Experiments
In this section, we conduct experiments on a real-world dataset from various cities to evaluate the performance of our KALM4REC framework. Specifically, we assess the candidate retrieval module and the LLMs’ capabilities in re-ranking restaurant candidates for cold-start user recommendations. Additionally, we report an extensive set of experiments to investigate the impact of various factors on the quality of the LLMs’ re-ranking.
4.1. Experimental Settings
Dataset. The experiments utilize the Yelp.com dataset (Vu et al., 2020), featuring over 67,000 restaurant reviews across three English-speaking cities. The dataset is partitioned into training and testing sets. Since we deal with the problem of cold-start user recommendation, we split the dataset such that users in the test set do not have any reviews in the train set. Dataset statistics are provided in Table 1.
Table 1. Data statistic of the Yelp dataset which contains reviews from three English-speaking cities. We split the train/test sets such that no users in the test set has review in the train set (cold-start user setting).
City # Restaurants # Reviews # Users
Train Test
Edinburgh 938 12,753 1,187 223
Singapore 983 12,878 1,046 197
London 986 42,061 4,396 825
Table 2. Performance comparison of various retrieval methods in Precision@20 and Recall@20 (in percentage). The highest performance is highlighted in bold, and the second highest performance is underlined.
City Conventional Ours Improve
Jaccard MF MVAE CBR MPG
P@20 R@20 P@20 R@20 P@20 R@20 P@20 R@20 P@20 R@20 P@20 R@20
[…]
Figure 4. Examine other impacts to KALM4REC . Left figure indicate effectiveness of keyword score, ”w” refers result for TF-IRF score mentioned in 6 while ”1” refers all scores are 1. Right figure indicate performance of different BERT model to encode the keyword for MPG to find the closest keyword in case of there is no matching keyword in training data. ”1” refers to all-MiniLM-L6-v2 while ”0” indicates whaleloops/phrase-bert. Both results are report in Precision@20 using MPG.
Keyword processing. We use Spacy [^2] https://spacy.io/ to extract keywords from reviews. Specifically, we keep consecutive words having targeted part-of-speech tags as keywords. The targeted part-of-speech tags include ‘ADJ’, ‘NOUN’, ‘PROPN’, and ‘VERB’ . These keywords encompass different word categories, describing the aspects mentioned in reviews. Across our study, keywords and reviews are vectorized by sBERT [^3] https://huggingface.co/sentence-transformers/all-MiniLM-L6-v2, a pretrained model to captures the semantic information.
Evaluation and implementation details. We evaluate our model using three metrics: recall (Recall@K), precision (Precision@K), and Normalized Discounted Cumulative Gain (NDCG@K). For the retrieval task, we focus on Recall@20 and Precision@20. To assess performance of re-ranking task across 20 candidates, various values of K, including 1, 3, and 5 are employed. These choices align with recent literature that focusing on the accurate of top rank items (He et al., 2023; Hou et al., 2024). Training, testing, and inference are conducted on Colab Pro with T4 GPU, using a batch size of 256 and the AdamW optimizer with a default learning rate of $1e^{-3}$. We provide the implementation code online for reproducibility. Additionally, experiments involving LLMs are conducted on Gemini Pro 1.5444https://deepmind.google/technologies/gemini/pro/, GPT-3.5-Turbo [^5] https://platform.openai.com/docs/models/gpt-3-5-turbo, Mistral 8x7B [^6] https://docs.mistral.ai/getting-started/models/, and LLama 3-8B [^7] https://huggingface.co/meta-llama/Meta-Llama-3-8B.
Retrieval Baselines. To evaluate our retrieval module, we compared it with the following retrieval baseline methods:
  • Jaccard similarity-based method measures the similarity between cold-start user and restaurants based on the overlapping keywords. The restaurants candidates are ranked based on Jaccard similarity scores.
  • Matrix Factorization (MF) is a classic collaborative filtering method that decomposes the user-item interaction matrix and use the BPR (Rendle et al., 2009) loss to optimized. The optimized function is described in Equation 3. After training, the model learns the relationships between the users in the training set and the restaurants. For cold-start user recommendation, we first retrieve similar users from the training set using Jaccard similarity based on the input keywords. Then, we rank the restaurants based on the average ranking scores of these similar users.
  • Multi-Variational Autoencoder (MVAE) is an item-based Collaborative Filtering method based on the variational autoencoder. It assumes that the data is generated from a multinomial distribution and uses variational inference for parameter estimation. We utilize the code provided by the authors [^8] https://github.com/younggyoseo/vae-cf-pytorch. For generating candidates in cold-start user recommendations, we employ the same strategy as with MF: the candidate scores are the average scores from similar training users, determined using Jaccard similarity based on input keywords.
Table 3. Performance in comparison with different retrieval models with Gemini. All ALMs obtain remarkable improvement against it backbone in retrieval module that indicate the effectiveness of our framework. P, R, N stands for Precision, Recall, and NDCG, respectively.
City Methods P@1 R@1 P@3 R@3 N@3 P@5 R@5 N@5
Edinburgh CBR 29.15 5.05 19.13 10.23 32.81 15.07 13.19 32.87
CBR-LLM 50.24 10.16 26.86 14.17 44.39 18.92 16.38 40.34
MPG 54.71 10.91 38.12 20.55 58.49 30.40 26.07 55.86
MPG-LLM 61.69 11.54 42.95 21.36 62.19 31.31 26.77 56.63
Singapore CBR 17.76 3.98 12.52 8.10 23.13 10.56 10.53 23.91
CBR-LLM 29.44 5.97 17.93 10.63 32.82 13.30 13.11 29.72
MPG 41.12 8.19 27.41 15.55 48.11 22.34 20.43 46.76
MPG-LLM 51.78 9.77 31.13 16.38 51.05 23.05 23.05 49.76
London CBR 26.06 5.04 17.45 9.59 31.46 14.52 13.09 32.87
CBR-LLM 38.30 7.75 22.66 12.94 39.30 17.31 16.06 39.34
MPG 47.76 9.60 34.34 19.47 52.46 27.68 25.17 52.26
MPG-LLM 55.15 11.19 36.48 20.48 55.40 28.53 26.89 54.03
4.2. Evaluating Candidates Retrieval
The goal of candidates retrieval is not only to achieve a high ranking (precision) but also to retrieve as many correct candidates as possible (recall). This ensures that the re-ranking module has a comprehensive set of candidates to work with in its input. Therefore, we evaluate the retrieval methods using precision and recall of the top 20 candidates, i.e., Precision@20 and Recall@20. The results are reported in Table 2. Column “Improve” indicates the improvement of our best retrieval model (MPG) against the best conventional approach. Table 2 shows that both MF and MVAE outperform Jaccard similarity-based method. However, these methods face challenges because they are trained on user-restaurant interactions rather than keywords, making it difficult to accurately capture the intentions of cold-start users as expressed through input keywords. In contrast, approaches leveraging keywords exhibit notable enhancements. CBR performs better than all the baselines method as it utilized keywords to compute user and restaurant representations. The best performance is achieved using MPG, which harnesses the graph structure between keywords and restaurants. MPG demonstrates consistent improvement across three cities, enhancing Recall@20 by at least 21.38% and Precision@20 by at least 4.75%. Furthermore, we examine the role of TF-IRF in constructing edges( Equation 6) and pretrained BERT model for the keyword vectorization. As shown in Figure 4, MPG with TF-IRF and ”all-MiniLM-L6-v2” outperforms other configurations.It is noteworthy that in our framework, achieving high recall is crucial, as these candidates will be re-ranked in the next phase.
4.3. Investigating the Re-ranking Capability of LLMs for Cold-start User Recommendations
In this section, we evaluate the re-ranking capability of LLMs within different settings and conditions. Specifically, we investigate the following research questions.
  • RQ1. Can LLMs re-rank candidates to improve the cold-start user recommendation task?
  • RQ2. Is employing keywords more effective than using reviews in prompting?
  • RQ3. How do prompt design strategies impact the efficiency of LLMs?
  • RQ4. The impact of keywords on the re-ranking performance.
  • RQ5. Do LLMs suffer from biases while ranking?
RQ1. Can LLMs re-rank candidates to improve the cold-start user recommendation task? We investigate the capability of using LLMs to re-rank candidates in cold-start user recommendation. Table 3 shows the comparison between retrieval only (CBR and MPG) and the full KALM4REC framework (i.e., retrieval and re-ranking using LLM – Gemini Pro). The results show the consistent improvements when using LLM-based candidates re-ranking for both CBR and MPG, for all the 3 cities regarding the evaluation metrics. Additionally, there is a significant improvement compared to sole retrieval modules, indicating the extensibility of our framework with alternative retrieval modules in the future.
Table 4. Performance of KALM4REC using different LLMs for re-ranking with MPG as the retrieval method. The results demonstrate that all LLMs enhance the precision and recall of the top recommended item, with Gemini consistently delivering the best overall performance. P, R, N stands for Precision, Recall, and NDCG metrics, respectively.
City Name Model P@1 R@1 P@3 R@3 N@3
Edinburgh MPG 54.71 10.91 38.12 20.55 58.49
[…]
Llama3 8B 56.50 11.27 36.62 18.91 56.42
Singapore MPG 41.12 8.19 27.41 15.55 48.11
[…]
Llama3 8B 45.18 8.71 29.10 14.99 49.27
London MPG 47.76 9.60 34.34 19.47 52.46
Gemini 55.15 11.19 36.48 20.48 55.40
[…]
Figure 5. Comparing the recommendation performance between the candidates retrieval module (MPG) and KALM4REC with the use of reviews or keywords. Results show that prompting using keywords obtain consistent improvements over MPG, whereas prompting using reviews only shows improvement at P@1 and R@1. Experiment is conducted using Gemini Pro.
Figure 6. Total number of prompt tokens when using reviews versus keywords to prompt Gemini Pro (in millions).
We also examine the performance of KALM4REC with different LLMs. Table 4 compares the performance of various LLM models –Gemini, GPT-3.5, Mistral 8B, and Llama3 8B– across multiple cities using precision at 1 (P@1), recall at 1 (R@1), precision at 3 (P@3), recall at 3 (R@3), and NDCG at 3 (N@3). We use MPG as the retrieval method. The results demonstrate that all LLMs enhance the precision and recall of the top recommended item (P@1 and R@1). Across all cities, Gemini consistently achieves the highest scores in almost every metric, demonstrating its robustness in identifying and ranking relevant restaurant candidates. For instance, Gemini attains the highest P@1 values, indicating superior accuracy in recommending the top restaurant. Additionally, Gemini’s strong performance in N@3 reflects its ability to rank relevant candidates effectively. GPT-3.5 and Mistral 8B models also demonstrate improvements over MPG. Both models show competitive results in P@3 and R@3, indicating their effectiveness in providing multiple relevant recommendations. GPT-3.5, for example, often comes close to Gemini’s performance, particularly in N@3, suggesting its capability to enhance the ranking quality significantly compared to MPG. The Mistral 8B model, while not always leading, maintains consistent and reliable performance across all metrics. Llama 3-8B, though slightly behind the other LLMs, still shows improvement over MPG in several metrics. While it may not be as effective in capturing the top preferences of users as the leading models, it provides a consistent alternative that surpasses the baseline.
RQ2. Is employing keywords more effective than using reviews in prompting? To convey user preferences and restaurant characteristics, the straightforward approach is to use the full content of user reviews. However, due to the context length limitations in LLM prompts, KALM4REC uses keywords instead. We compare the effectiveness of LLMs when using reviews versus keywords for the re-ranking task. For each restaurant candidate in a prompt, we describe it using a representative review, which is selected based on its distance to the review centroid. The centroid embedding is calculated as the average of the embeddings of all the restaurant’s reviews. We use Gemini Pro for this evaluation. The comparison results across 3 cities are illustrated in Figure 5 in Precision@k and Recall@k with k = 1, 3. From Figure 5, we find that rankings by our KALM4REC consistently show remarkable improvements across all cities in P@1 and R@1. However, we observe KALM4REC with review gets lower number in P@3 and R@3 compared to the retrieval model for Singapore. Whereas, using keywords offers improvements across all the metrics. This improvement is likely attributed to the detailed information provided by keywords, which directly correspond to the input keywords. This enables the LLM to better understand the relevance between the input user and the candidate restaurants.
Another benefit of using keywords is the length-efficiency when prompting LLMs. Figure 6 depicts the total number of prompt tokens used when utilizing reviews versus keywords to prompt Gemini Pro. Using reviews required approximately 800,000 tokens for Singapore, 950,000 for Edinburgh, and 3,190,000 for London. In contrast, using keywords required only about 170,000, 340,000, and 718,000 tokens for the respective cities. In total, prompts using reviews are four times longer than those using keywords. Keywords not only outperform reviews in effectiveness but are also much more efficient in terms of prompt length and associated costs.
RQ3. How do prompt design strategies impact the efficiency of LLMs? We compare the performance of KALM4REC with zero-shot [$\mathcal{H}_{z}$] and few-shot [$\mathcal{H}_{f}$] strategies. In this experiment, we set number of keywords for user ranging from 3 to 5 while number of keywords for restaurant range from 3 to 20. We report the best setting for each strategy in Figure 7. The experiments is conducted with Gemini Pro. We observe significant performance improvements across all strategies of zero-shot and few-shot when compared to MPG. The number of shots (1 shot, 2 shots, 3 shots) indicates the number of examples introduced to LLMs. Overall, increasing the number of examples enhances LLM performance, with the 3-shot approach mostly achieve the best results. This outcome is reasonable as LLMs can better capture user intent with more examples.
RQ4. The impact of keywords on the re-ranking performance. We investigate the impact of keywords on the re-ranking performance regarding the number of keywords and their orders. Specifically, we compare the re-ranking results when varying the number of keywords used to describe restaurants and users. We also examine whether the order of keywords affect the performance of KALM4REC. Figure 8 shows the Recall@1 of zero-shot and one-shot for Singapore using Gemini Pro. We observed that one-shot requires fewer keywords yet still achieves relatively good performance. Figure 9 shows the evaluation performed on Edinburgh and Singapore using Gemini Pro when we change number of keywords and whether to shuffle the order of the keywords in prompts. The results show that the number of keywords used for restaurants influences LLM performance. Specifically, in Edinburgh, LLM performance increases with the number of keywords (with order). The visualizations also emphasize the crucial role of keyword order in the prompt. Performance notably declines when keywords are not ordered according to our ranking, indicating that LLMs better comprehend user preferences when keywords are sorted by their importance score. This effect is particularly pronounced in Edinburgh, where performance without the ordered keywords is even lower than that of MPG.
Figure 7. The impact of different ranking strategies on performance across various cities. The red dashed line represents results from the retrieval model only. Experiement is conducted on Gemini.
Figure 8. Comparison the number of keywords used to describe restaurant candidates and users in zero-shot and one-shot scenarios in Recall@1, we find that one-shot requires fewer keywords yet still achieves relatively good performance. Experiment is conducted using Gemini Pro on Singapore data.
(a) Edinburgh
(b) Singapore
Figure 9. Performance comparison in terms of the number of keywords and their order of representation for each restaurant within the zero-shot design. Experiment is conduct with Gemini for Edinburgh.
RQ5. Do LLMs suffer from biases while ranking?
We adopt the insights from (Hou et al., 2024) that the order of candidates affects the ranking results of LLMs. As the result, we make an experiment where the order of restaurant candidates is shuffled. The experiment is conducted with Gemini Pro and depicted in Figure 10. The shown results indicate that Gemini are generally sensitive to the order of examples in the prompt. Especially, the ranking performance by shuffle drops significantly for Singapore and London, lower than performance of retrieval models. This suggests that LLMs perform well when integrated with efficient retrieval models, as demonstrated in our previous experiment (Table 4), highlighting the crucial role of our proposed retrieval model.
Figure 10. Impact of candidate order in prompt to our KALM4REC. Experiment is conducted using Gemini Pro.
5. Discussion and Conclusion
In this work, we investigate the capabilities of Augmented LLMs for cold-start user recommendations by leveraging keywords extracted from user reviews. Our experimental results demonstrate that the candidates retrieval and LLM-based candidates ranking methods proposed in KALM4REC effectively address cold-start issues. For retrieving potential restaurant candidates, we present CBR and MPG, keyword-based methods, which outperform conventional approaches. We then employ LLMs to re-rank the obtained candidates using designed prompting strategies that incorporate keywords to represent users and restaurants. Comprehensive experiments indicate that KALM4REC is capable of handling cold-start user scenarios. Additionally, our framework shows the potential of integrating different retrieval and language models to achieve promising performance in the future.
We acknowledge that certain factors may affect the performance of our work. For instance, while the constructed graph in MPG is effective, the connections between keyword nodes should also be considered, as similar keywords may need to pass information to each other before reaching the restaurant nodes. However, implementing this could be computationally costly due to the vast number of keywords in large cities like London. Additionally, the number of keywords introduced into the prompt may be reduced by increasing the number of example introductions, as shown in the experiments. For future work, we plan to develop technical approaches to enhance the robustness of each stage in our framework, leveraging the insights drawn from our comprehensive set of experiments.