### v2

# Keyword-driven Retrieval-Augmented Large Language Models for Cold-start User Recommendations

Abstract.
Recent advancements in Large Language Models (LLMs) have shown significant potential in enhancing recommender systems. However, addressing the cold-start recommendation problem, where users lack historical data, remains a considerable challenge. In this paper, we introduce KALM4Rec (Keyword-driven Retrieval-Augmented Large Language Models for Cold-start User Recommendations), a novel framework specifically designed to tackle this problem by requiring only a few input keywords from users in a practical scenario of cold-start user recommendations. KALM4Rec operates in two main stages: candidates retrieval and LLM-based candidates re-ranking. In the first stage, keyword-driven retrieval models are used to identify potential candidates, addressing LLMs’ limitations in processing extensive tokens and reducing the risk of generating misleading information. In the second stage, we employ LLMs with various prompting strategies, including zero-shot and few-shot techniques, to re-rank these candidates by integrating multiple examples directly into the LLM prompts. Our extensive evaluation, using Yelp restaurant data from three English-speaking cities and TripAdvisor hotel data, demonstrates that KALM4Rec excels in improving recommendation quality across the two domains. The framework’s adaptability to different domains highlights its potential for widespread applications. By integrating in-context instructions with LLMs, KALM4Rec notably enhances the performance of cold-start recommender systems, offering a novel approach to exploring solutions in this field. Our code is available at …
Restaurant Recommendation, Cold-start User Recommendation, Large Language Model, Prompt Tunning.
isbn: 978-1-4503-XXXX-X/18/06ccs: Information systems Retrieval models and rankingccs: Information systems Language modelsccs: Information systems Learning to rank
1. Introduction
Recommender systems play a pivotal role in assisting users in navigating the vast number of available choices in the digital world. However, a significant and ongoing challenge in this field is addressing the issue of cold-start users. These are users who are new to the platform and therefore have little to no interaction history. The lack of data makes it difficult for the system to generate accurate and personalized recommendations. Conventional Collaborative Filtering (CF) such as NCF(He et al., 2017), BPR (Rendle et al., 2009)and LightGCN (He et al., 2020) or generative methods (Liang et al., 2018; Wang et al., 2023c; Liu et al., 2023b) struggle to suggest relevant items effectively for these new users due to the lack of detailed preference information. While user-user content-based algorithms offer a solution by relying on user features to find similar users and recommend positively interacted items, this approach raises privacy concerns. Overall, there remains a significant gap in research addressing the needs of cold-start users, highlighting the necessity for further exploration in this area.
Figure 1. A pictorial illustration of our Keyword-driven Retrieval-Augmented Large Language Models for cold-start users recommendation (KALM4REC).
Meanwhile, Large Language Models (LLMs) have garnered considerable interest for their potential to enhance recommender systems, showcasing significant proficiency in natural language understanding, generation, and complex reasoning. Leveraging their robust architecture and extensive training on vast datasets, several studies employ LLMs to comprehend and address user requirements throughout the recommendation process (Gao et al., 2023; Wang et al., 2023b; Li et al., 2024; Wu et al., 2023). Moreover, researchers have explored LLMs’ potential to alleviate the cold-start recommendation problem through in-context learning (He et al., 2023). With LLMs’ adaptability to new tasks, prompts are often tailored to incorporate user and item information, facilitating effective adaptation and performance enhancement (Sanner et al., 2023; Wang et al., 2024b).
Challenge. Despite their excellent capacities, existing LLMs suffer from several limitations when applied to recommender systems, especially for cold-start scenarios. For one, lacking comprehensive knowledge in a specific domain may result in non-factual outputs. While fine-tuning smaller LLMs could potentially reduce such hallucinations, this solution is often impractical due to the significant resources required (Mialon et al., 2023). Furthermore, directly incorporating user/item information can be costly in terms of token usage and restricts input length due to limited context length.
Approach. For recommender systems in the various domains such as restaurant, user reviews offer valuable insights into their opinions about businesses. Modeling these personal reviews is beneficial for comprehending users’ preferences and capturing the characteristics of restaurants (Vu et al., 2020). For cold-start users who lack historical data, a practical strategy is to ask them to provide a few keywords describing their preferences. This approach helps create an initial user profile with minimal effort and without compromising their privacy. Furthermore, using keywords instead of entire reviews to prompt LLMs can significantly enhance the efficiency and accuracy of the recommendations while minimizing token usage and computational costs.
To this end, we present a novel framework called Keyword-Driven Retrieval-Augmented Large Language Models for Cold-Start User Recommendations (KALM4REC). Our framework is built upon two essential components centered around keywords: candidates retrieval focusing on retrieving relevant items, and LLM-ranker which leverages LLM to re-rank the retrieved candidates (illustrated in Figure 1). User reviews reflect user preferences but often contain noise, potentially leading to inaccurate representations of user preferences. By exclusively leveraging extracted keywords, we retain the contextual essence of the reviews while still describing the user profile well. Utilizing keywords also allows us to optimize the prompts by reducing noise, thereby enhancing their effectiveness while still ensuring sufficient information of users and restaurants in LLMs, effectively addressing cold-start issues. Essentially, KALM4REC, grounded in sets of keywords, aims to narrow down the candidate set relevant to predefined cold-start preferences. The LLM utilizes its contextual understanding and reasoning capabilities to generate a ranked list of items for recommendations. Our contributions can be summarized as follows:
  • To the best of our knowledge, KALM4REC stands as the pioneering framework to employ keyword as the focal point of both the candidate retrieval and LLM-based re-ranking modules in recommender system. By prioritizing keywords, our augmented LLMs framework effectively addresses the significant challenge of hallucinated outcome and optimizing prompt design, resulting in promising performance for cold-start users recommendation.
  • We propose approaches that take advantages of the extracted keywords to obtain most suitable candidates for cold-start users. Furthermore, we designed prompting templates to enable LLMs work efficiently for re-ranking tasks.
  • Finally, we conduct comprehensive experiments to validate and demonstrate the effectiveness of KALM4REC, compared to previous baseline methods for the cold-start user recommendation challenge, using two different domains including restaurant (Yelp) and hotel (Tripadvisor).
2. Related works
Cold-start Recommendations. A recommendation system that aims to provide suggestions to new users with little or no historical data is known as a cold-start recommendation system. These systems play a critical role in businesses and services that require recommendations for new users or products. Various strategies have been proposed to tackle the challenge of generating accurate recommendations in such scenarios (Guo et al., 2020; Bianchi et al., 2017; Gao et al., 2018; Camacho and Alves-Souza, 2018). Collaborative filtering (CF) (He et al., 2017; Wang et al., 2019a) has traditionally been effective in recommendation systems but struggles with the cold-start problem. To address this challenge, some approaches have integrated data augmentation techniques into CF models, incorporating user and item information to improve recommendations (Bianchi et al., 2017; Gao et al., 2018). While methods such as (Wei et al., 2020; Anwar et al., 2022) extend CF techniques, others like (Son and Kim, 2017) combine collaborative filtering with content-based methods to mitigate the cold-start issue. Deep learning techniques have also been employed to create representations or embeddings that capture the latent features of users and items, aiding in cold-start scenarios (Ying et al., 2018; Volkovs et al., 2017; Zhao et al., 2022; Wang et al., 2019b). DropoutNet (Volkovs et al., 2017) directly addresses cold-start issues by avoiding the need to aggregate all content-based information into the objective function. (Zhao et al., 2022) employs latent variables to develop a probabilistic approach for learning distributions over side information, helping to ”warm up” cold item embeddings. Additionally, graph-based techniques (Ying et al., 2018) leverage user-item interaction graphs to infer user preferences, while (Wang et al., 2019b) uses knowledge graphs as a source of side information to support recommendations. Some methods generate warm embeddings for cold-start items based on their features and ID embeddings, utilizing meta-networks (Zhu et al., 2021). Moreover, (Zhou et al., 2023) explores contrastive learning with CF-based modules to create content-based collaborative embeddings using information from warm-start items to address the cold-start issue. Although these methods show improvements in cold-start problems, they rely heavily on auxiliary information which usually causes privacy concerns.
LLMs for Recommendations.
Figure 2. KALM4REC begins with the Keyword Processing module, where we extract noun phrases as keywords from user reviews to construct their profiles. For cold-start users, candidate restaurants are generated using a retrieval model based on their selected terms. These candidates are then passed to the Prompt Generation stage, where information about user preferences and restaurant characteristics is introduced. Finally, leveraging their pre-trained knowledge, LLMs capture user expectations and perform re-ranking accordingly.
The impressive capabilities of Large Language Models (LLMs) in natural language understanding, generation, and complex reasoning have inspired innovative approaches in recommendation systems (Fan et al., 2023; Geng et al., 2022; Wang et al., 2024a, 2023b; Li et al., 2024; Wu et al., 2023; Dai et al., 2023; Wang et al., 2023a). Among these, the ”Pretrain, Personalized Prompt, and Predict Paradigm” (P5) (Geng et al., 2022; Xu et al., 2024) stands out as a groundbreaking approach, harnessing the inherent flexibility of the language to tackle diverse recommendation tasks with remarkable success. Rather than fine-tuning models (Lyu et al., 2024; Geng et al., 2022; Cui et al., 2022), some methods use LLMs in a zero-shot setting to recommend items directly, demonstrating promising results (Liu et al., 2023a; Zhang et al., 2023; Kang et al., 2023; Wang and Lim, 2023; Sun et al., 2024). To address the issue of hallucinations when being used for direct recommend ations, LLMs can be employed as re-rankers within recommendation systems(Hou et al., 2024; Dai et al., 2023). For example, Dai et al. (2023) found that GPT could outperform weak baselines when re-ranking a list of candidate items. A notable example of LLMs’ potential is highlighted by Hou et al. (2024), which used LLMs as zero-shot rankers. They transformed the recommendation process into a conditional ranking challenge based on sequential interaction histories and carefully crafted prompt templates. This approach not only demonstrates LLMs’ ability to deliver contextually accurate recommendations but also indicate that by enhancing retrieval modules, LLMs can provide a richer and more responsive user experience. Furthermore, Sanner et al. (2023); Wang et al. (2024b) explore the integration of user and item information into prompts to overcome cold-start issues, while He et al. (2023) shows that in-context learning with LLMs can effectively address these challenges without needing extra data. Despite these advances, there remains a significant gap: no research has yet investigated the combination of keyword extraction with LLMs to boost recommendations for cold-start users. Our study aims to bridge this gap by introducing a novel approach that blends keyword extraction with LLMs to enhance recommendation accuracy in cold-start scenarios.
3. Proposed Framework
In this section, we present the details of our keyword-driven framework, KALM4REC (Figure 2), designed to tackle the cold-start recommendation problem. KALM4REC unfolds in two primary stages: (1) Candidates retrieval, where meaningful noun phrases are extracted from collected reviews to form word sets representing users and targeted items, and subsequently utilized for queries from cold-start users to obtain potential candidates. This phase employs our retrieval models, leveraging a collaborative filtering mechanism with keywords at its core to rank potential candidates; and (2) LLM-based candidates re-ranking, where an LLM is employed as a ranker to re-rank the retrieved candidates.
3.1. Problem Formulation
Given the set of users $\mathcal{U}$ and the set of targeted items (items for short) $\mathcal{R}$, the interaction between a user $u\in\mathcal{U}$ and an item $r\in\mathcal{R}$ is presented as a set of reviews =$\mathcal{I}=\{i_{u,r}\}$, where each entry $i_{u,r}$ represents a review if user $u$ has a review for item $r$. The task involves recommending relevant items to a new user $u_{c}$ (cold-start user). The new user can opt to declare their preference via a set of pre-defined keywords $k_{u_{c}}$. Simultaneously, user reviews also contain “keywords” $\mathcal{K}=\{k_{u,r}\}$, where each entry $k_{u,r}$ is a set of keywords extracted from $i_{u,r}$. We denote $k_{u}$ as a set of keywords extracted from all user $u$’s reviews, and $k_{r}$ as a set of keywords extracted from all reviews written for item $r$. The problem of keyword-driven cold-start user recommendation is defined as follows. Given a set of keywords $k_{u_{c}}$ from a cold-start (new) user, return a ranked list of relevant items $R_{u_{c}}$ where $R_{u_{c}}\subset\mathcal{R}$. In the following, we now present KALM4REC with the details of the two stages.
3.2. Candidates Retrieval
In this stage, we present methodologies for retrieving potential candidates that include Content-based Recommendation by Keywords (CBR) and Message Passing on Graph (MPG). For CBR, we utilize deep neural networks to learn the interaction function from sets of keywords representing users and items. To accommodate the arbitrary order of keywords within sets, we adopt a pooling layer based on an attention mechanism. For MPG, we construct a heterogeneous graph comprising nodes representing keywords and items, with two types of edges defined, and introduce a scheme for estimating scores of the connections between keywords and items.
3.2.1. Content-based Recommendation by Keywords (CBR)
Motivated by the effectiveness of existing works (Chen et al., 2022), we generate the embedding for the user and items from its keyword representation. For CBR, profiles of all users including cold-start users is presented as sequences of keywords.
Given the sequence of keywords for user $u$: $E_{u}=[E_{u}^{1},E_{u}^{2},...,E_{u}^{T}]\in\mathbb{R}^{T\times d}$, where $T$ denotes the number of keywords for the user and $d$ is the dimension size of the keyword vectorized from sBERT (Reimers and Gurevych, 2019). The recommendation problem can be abstracted as $\hat{y}_{ur}=f(u,r|\Theta)$, where $f$ is a function learned to estimate the score for the user-item pair. Let $p_{u}$ and $q_{r}$ denote the latent vectors for user $u$ and item $r$, learned by $\Theta$. The interaction can be estimated as the inner product of $p_{u}$ and $q_{r}$:
(1) $\hat{y}_{ur}=f(u,r|p_{u},q_{r})=p_{u}^{T}q_{r}$
where $p_{u}$ and $q_{r}$ are learned to encode the user profile and restaurant profile, respectively, through a set pooling mechanism based on an attention mechanism (Safari et al., 2020) denoted as $G(E_{u})$.
(2) $G(E_{u})=softmax(W_{c}E_{u}^{T})E_{u}$
Here we use a softmax layer, and $W_{c}\in\mathbb{R}^{d}$ inferring a trainable parameter. $G(E_{u})$ can be interpreted as a dot product attention, where the keys and the values correspond to the same representation, and the query is only a trainable parameter. Therefore, $G(E_{u})$ represents a weighted average of the encoder sequence of features. The learned features are expected to fully represent the user profile.
For optimization, we adopt Bayesian Personalized Ranking (BPR) (Rendle et al., 2009), where $r_{p}$ is the item that user $u$ prefers over $r_{n}$. The model is optimized by:
(3) $\sum_{(u,r_{p},r_{n})\in\mathcal{D}_{S}}\ln\sigma(\hat{y}_{u,r_{p},r_{n}})-\lambda_{\Theta}\|\Theta\|^{2}$
where $\mathcal{D}_{S}=\{(u,r_{p},r_{n})|r_{p}\in\mathcal{I}^{+}_{u}\land j\in\mathcal{I}^{-}_{u}\}$ is the training set, $\mathcal{I}^{+}_{u}$ denoting the items the user $u$ liked and $\mathcal{I}^{-}_{u}$ indicating some items user $u$ doesn’t like; $\hat{y}_{u,r_{p},r_{n}}$ represents the relation between user $u$ and items $r_{p}$ and $r_{n}$; $\sigma$ is the logistic sigmoid function; $\lambda_{\Theta}$ are model regularization parameters.
For inference, the process is similar to training, as both stages involve encoding sets of keywords.
3.2.2. Message Passing on Graph (MPG)
Utilizing CBR alone is not sufficient to provide satisfactory embeddings for Collaborative Filtering. The primary deficiency lies in the embedding function’s failure to encode crucial collaborative signals latent in keyword-item and keyword-user interactions. Drawing inspiration from LightGCN (He et al., 2020), which leverages structured information, we propose a model based on a tripartite graph structure (user - keyword - item). Specifically, we construct the graph using training data, with edges established for (user - keyword) and (keyword - item) interactions when a user uses a keyword to review and a item’s review contains that keyword. Information is passed through the edges in the graph, and node information could be generated as follows:
(4) $q_{r}=AGG(q_{w},w\in k_{r})$
where $AGG$ is a function that aggregates information from neighboring nodes, and similarly for $q_{w}$.
However, due to the challenge posed by a large number of nodes, reaching millions, we propose an unsupervised learning model. First, we adopt an aggregator $(AGG)$ based on a simple weighted sum (similar to the work from LightGCN (He et al., 2020)), but without trainable parameters to determine the weights.
Figure 3. Graph construction with connections between users, keywords, and items. Green nodes represent keywords from the training data, while dashed lines indicate connections established via nearest neighbors for keywords, not in the data. Edge brightness reflects the importance score.
For an item, the message is generated by multiplying the information of the keyword (source node) by its edge weight, then summing up the messages to update as follows:
(5) $q_{r}=\sum_{r\in\mathcal{R},w\in k_{r}}e^{w}*a^{w}_{r}$
where $e^{w}$ denotes feature of keyword nodes $w$; $a^{w}_{r}$ denote the connection between node $w$ and node $r$. Specifically, the edge between a keyword node and an item can be represented by a value indicating the importance of a keyword to the item, denoted as $\mathcal{A}_{r}^{w}$. We introduce a scheme named TF-IRF to measure the importance score of a keyword to an item, which is similar to the TF-IDF idea. The score assigned to the edge established by an item $r$ and a keyword $w$ is denoted as $a^{w}{r}$ and defined as:
(6) $a_{r}^{w}=tf_{r}^{w}*irf^{w}$
with $tf_{r}^{w}=\frac{f_{r}^{w}}{q_{w}}$, where $f_{r}^{w}$ represents the number of times the term $w$ appears in the set of keywords for item $r$, and $q_{w}$ indicates the total number of times the term $w$ is used in the entire training data; $irf^{w}=\log\left(\frac{|\mathcal{R}|}{f^{w}}\right)$, where $f^{w}$ denotes the number of items containing the term $w$. The scores of all items $\mathcal{S}$ can be rewritten in matrix form as:
(7) $\mathcal{S}=\mathcal{M}_{w}\times\mathcal{A}_{r}^{w}$
where $\mathcal{S}\in\mathbb{R}^{1\times r},\mathcal{A}_{r}^{w}\in\mathbb{R}^{w\times r}$, and $\mathcal{M}_{w}\in\mathbb{R}^{1\times w}$ denotes the occurence matrix represents the connection between cold-start user and their selected keywords. Then, the top-k possible items are provided by:
(8) $C_{k}=\operatorname*{argmax}_{k=20}(\mathcal{S})$
Notably, no training is needed for this approach, only graph construction. If cold-start users select keywords absent in the training data, we replace them with the most similar keywords in semantics from the training data. To achieve this, we employ a pretrained BERT model (Reimers and Gurevych, 2019) to vectorize all keywords, ensuring that semantically similar keywords are close to each other. The most similar keyword then searched by using nearest neighbor model [^1] https://scikit-learn.org/stable/modules/neighbors.html.
3.3. LLM-based Candidates Re-ranking
[…]
Input: Please suggest the 15 most suitable restaurants for me from the candidate set that I will visit them, according to the user and candidate restaurant keywords I provided above.
Output: Must include 15 restaurants in the candidate restaurant set, not keyword. Desired format is string: restaurant_id_1, restaurant_id_2, …
Following, we propose a general item recommendation prompt framework using keywords, consisting of (1) user keywords; (2) candidate sets; and (3) item keyword sets. This prompt framework is denoted as [$\mathcal{H}_{z}$] and [$\mathcal{H}_{f}$] represent for zero-shot and few-shot pattern, respectively.
3.3.1. Zero-shot $\mathcal{H}_{z}$
[…]
The candidate restaurant set for me is enclosed in square brackets, with the restaurants separated by commas (Format: [restaurant_id_1, restaurant_id_2,…]) is: {candidate restaurant set}
Keywords associated with candidate restaurants have the following form: restaurant_id_1 (keyword 1, keyword 2,…) are {restaurant keywords sets}.
3.3.2. Few-shot [$\mathcal{H}_{f}$]
Similar to human reasoning, LLMs benefit from good examples to better understand human intentions and criteria for desired answers. Therefore, few-shot learning often yields better performance than zero-shot. The pattern of few-shot prompts [$\mathcal{H}_{f}$] above comprises examples and the information in [$\mathcal{H}_{z}$] then filled into Template $\mathcal{T}$. The keywords of candidates are ordered in our prompts to preserve the order from retrieval models, which is based on TF-IRF scores.
[…]
# Repeat $\textbf{i}\in\{1,..,k\}$:
Example $i$:
There are the keywords that user often mention when wanting to choose restaurants: {user keywords}.
The candidate restaurant set for user is enclosed in square brackets, with the restaurants separated by commas (Format: [restaurant_id_1, restaurant_id_2,…]) is: {candidate restaurant set}
Keywords associated with candidate restaurants have the following form: restaurant_id_1 (keyword 1, keyword 2,…) are {restaurant keywords sets}.
You should rank a list of recommendations for user as follows: {restaurant_id_1, restaurant_id_2, …}
Pattern: [ $\mathcal{H}_{z}$ ]
4. Experiments
In this section, we conduct experiments on a real-world dataset from various cities to evaluate the performance of our KALM4REC framework. Specifically, we assess the candidate retrieval module and the LLMs’ capabilities in re-ranking candidates for cold-start user recommendations. Additionally, we report an extensive set of experiments to investigate the impact of various factors on the quality of the LLMs’ re-ranking.
4.1. Experimental Settings
Dataset. The experiments utilize two datasets: the Yelp.com dataset (Vu et al., 2020) and the TripAdvisor dataset (Li et al., 2013). The Yelp.com dataset features over 67,000 restaurant reviews across three English-speaking cities. The TripAdvisor dataset consists of 878,561 reviews from 4333 hotels crawled from TripAdvisor.com across 17 cities. The dataset is partitioned into training and testing sets. Since we deal with the problem of cold-start user recommendation, we split the dataset so that users in the test set do not have any reviews in the train set. Dataset statistics are provided in Table 1.
Table 1. Statistic of the datasets. In the Yelp dataset, ”Items” refers to restaurants, while in the TripAdvisor dataset, ”Items” refers to hotels.”
Dataset City # Items # Reviews # Users
Train Test
Yelp Edinburgh 938 12,753 1,187 223
Singapore 983 12,878 1,046 197
London 986 42,061 4,396 825
TripAdvisors 4333 878,561 9,579 1,900
Table 2. Performance comparison of various retrieval methods in Precision@20 and Recall@20 (in percentage). The highest performance is highlighted in bold, and the second highest performance is underlined.
Dataset Conventional Our Retrievals Improve
Jaccard MF MVAE CBR MPG
P@20 R@20 P@20 R@20 P@20 R@20 P@20 R@20 P@20 R@20 P@20 R@20
[…]
Figure 4. The left figure indicates the effectiveness of the keyword score, ”weighted” refers result for TF-IRF score mentioned in Equation 6 while ”all_ones” refers to assigning all edge’s scores to 1. The right figure indicates the performance of different pretrained BERT models to vectorized the keywords. ”MiniLM” refers to all-MiniLM-L6-v2 while ”whaleloops” indicates whaleloops/phrase-bert.
Keyword processing. We use Spacy [^2] https://spacy.io/ to extract keywords from reviews. Specifically, we keep consecutive words having targeted part-of-speech tags as keywords. The targeted part-of-speech tags include ‘ADJ’, ‘NOUN’, ‘PROPN’, and ‘VERB’ . These keywords encompass different word categories, describing the aspects mentioned in reviews. Across our study, keywords and reviews are vectorized by sBERT [^3] https://huggingface.co/sentence-transformers/all-MiniLM-L6-v2, a pretrained model to captures the semantic information.
Evaluation and implementation details. We evaluate our model using three metrics: recall (Recall@K), precision (Precision@K), and Normalized Discounted Cumulative Gain (NDCG@K). For the retrieval task, we focus on Recall@20 and Precision@20. To assess performance of re-ranking task across 20 candidates, various values of K, including 1, 3, and 5 are employed. These choices align with recent literature that focusing on the accurate of top rank items (He et al., 2023; Hou et al., 2024). Training, testing, and inference are conducted on Colab Pro with T4 GPU, using a batch size of 256 and the AdamW optimizer with a default learning rate of $1e^{-3}$. We provide the implementation code online for reproducibility. Additionally, experiments involving LLMs are conducted on Gemini Pro 1.5444https://deepmind.google/technologies/gemini/pro/, GPT-3.5-Turbo [^5] https://platform.openai.com/docs/models/gpt-3-5-turbo, Mistral 8x7B [^6] https://docs.mistral.ai/getting-started/models/, and LLama 3-8B [^7] https://huggingface.co/meta-llama/Meta-Llama-3-8B.
Retrieval Baselines. Retrieval Baselines. To evaluate our retrieval module, we compared it with the following retrieval baseline methods:
  • Jaccard similarity-based method measures the similarity between cold-start users and items based on the overlapping keywords. The candidates are ranked based on Jaccard similarity scores.
  • Matrix Factorization (MF) is a classic collaborative filtering method that decomposes the user-item interaction matrix and uses the BPR (Rendle et al., 2009) loss to optimize.
  • Multi-Variational Autoencoder (MVAE) is an item-based Collaborative Filtering method based on the variational autoencoder. We utilize the code provided by the authors [^8] https://github.com/younggyoseo/vae-cf-pytorch.
For cold-start user recommendation in MF and MVAE, we retrieve similar users from the training set using Jaccard similarity based on the input keywords. Then we estimate the item score for cold-start users based on the average ranking scores of their similar users.
Table 3. Performance in comparison with different retrieval models with Gemini. All LLMs obtain remarkable improvement against their backbone in the retrieval module.
Dataset Methods P@1 R@1 P@3 R@3 N@3 P@5 R@5 N@5
Yelp-Edinburgh CBR 29.15 5.05 19.13 10.23 32.81 15.07 13.19 32.87
CBR-LLM 50.24 10.16 26.86 14.17 44.39 18.92 16.38 40.34
MPG 54.71 10.91 38.12 20.55 58.49 30.40 26.07 55.86
MPG-LLM 61.69 11.54 42.95 21.36 62.19 31.31 26.77 56.63
Yelp-Singapore CBR 17.76 3.98 12.52 8.10 23.13 10.56 10.53 23.91
CBR-LLM 29.44 5.97 17.93 10.63 32.82 13.30 13.11 29.72
MPG 41.12 8.19 27.41 15.55 48.11 22.34 20.43 46.76
MPG-LLM 51.78 9.77 31.13 16.38 51.05 23.05 23.05 49.76
Yelp-London CBR 26.06 5.04 17.45 9.59 31.46 14.52 13.09 32.87
CBR-LLM 38.30 7.75 22.66 12.94 39.30 17.31 16.06 39.34
MPG 47.76 9.60 34.34 19.47 52.46 27.68 25.17 52.26
MPG-LLM 58.06 11.77 37.09 20.67 56.85 29.18 26.24 54.30
TripAdvisor CBR 21.38 3.35 15.30 7.16 28.92 12.99 10.00 31.23
CBR-LLM 36.27 5.79 20.05 9.42 36.69 14.84 11.49 11.49
MPG 54.60 8.73 41.92 19.72 61.36 33.52 25.98 58.99
MPG-LLM 59.43 9.54 42.15 19.87 62.36 33.58 25.98 59.65
Figure 5. Comparing the performance between the candidates retrieval module (MPG) and KALM4REC with the use of reviews or keywords. Results reported at Precision@1 and Recall@1. The experiment is conducted using Gemini Pro.
4.2. Evaluating Candidates Retrieval
The goal of candidates retrieval is not only to achieve a high ranking (precision) but also to retrieve as many correct candidates as possible (recall). This ensures that the re-ranking module has a comprehensive set of candidates to work with in its input. Therefore, we evaluate the retrieval methods based on P@20 and R@20, which measure the precision and recall of the top 20 candidates, respectively. The results are presented in Table 2, where the column ”Improve” highlights the performance improvement of our best retrieval model, MPG, compared to the best conventional method. Our findings reveal that both MF and MVAE outperform Jaccard similarity-based method. However, these methods face challenges because they are trained on user-item interactions rather than keywords, making it difficult to accurately capture the intentions of cold-start users as expressed through input keywords. In contrast, approaches leveraging keywords exhibit notable enhancements. CBR performs better than all the baselines method as it utilized keywords to compute user and item representations. The best performance is achieved using MPG, which harnesses the graph structure between keywords and items. MPG demonstrates consistent improvement across three cities, enhancing Recall@20 by at least 21.38% and Precision@20 by at least 4.75%. Furthermore, we examine the role of TF-IRF in constructing edges (Equation 6) and pretrained BERT models for the keyword vectorization. As shown in Figure 4, MPG with TF-IRF and ”all-MiniLM-L6-v2” outperforms other configurations. It is noteworthy that in our framework, achieving high recall is crucial, as these candidates will be re-ranked in the next phase.
4.3. Investigating the Re-ranking Capability of LLMs for Cold-start User Recommendations
In this section, we evaluate the re-ranking capability of LLMs within different settings and conditions. Specifically, we investigate the following research questions.
RQ1. Can LLMs re-rank candidates to improve the cold-start user recommendation task? We investigate the capability of using LLMs to re-rank candidates in cold-start user recommendation. Table 3 shows the comparison between retrieval only (CBR and MPG) and the full KALM4REC framework (i.e., retrieval and re-ranking using LLM – Gemini Pro). The results show consistent improvements when using LLM-based candidates re-ranking for both CBR and MPG, for all the datasets regarding the evaluation metrics. Additionally, there is a significant improvement compared to sole retrieval modules, indicating the extensibility of our framework with alternative retrieval modules in the future.
Table 4. Performance of KALM4REC using different LLMs for re-ranking with MPG as the retrieval method. The results demonstrate that all LLMs enhance the precision and recall of the top recommended item, with Gemini consistently delivering the best overall performance.
City Name Model P@1 R@1 P@3 R@3 N@3
Edinburgh MPG 54.71 10.91 38.12 20.55 58.49
[…]
Llama3 8B 56.50 11.27 36.62 18.91 56.42
Singapore MPG 41.12 8.19 27.41 15.55 48.11
[…]
Llama3 8B 45.18 8.71 29.10 14.99 49.27
London MPG 47.76 9.60 34.34 19.47 52.46
Gemini 58.06 11.77 37.09 20.67 56.85
[…]
(a) Number of token
(b) Fewshot performance
Figure 6. Left figure: Total number of prompt tokens when using reviews versus keywords to prompt Gemini Pro (in millions). Right figure: The impact of different ranking strategies on performance across various cities. The dashed line represents results from the retrieval model only.
We also examine the performance of KALM4REC with different LLMs. Table 4 compares the performance of various LLM models –Gemini, GPT-3.5, Mistral 8B, and Llama3 8B– across multiple cities in Yelp dataset using Precision at 1 (P@1), Recall at 1 (R@1), Precision at 3 (P@3), Recall at 3 (R@3), and NDCG at 3 (N@3). We use MPG as the retrieval method. The results demonstrate that all LLMs enhance the precision and recall of the top recommended item. Across all cities, Gemini consistently achieves the highest scores in almost every metric, demonstrating its robustness in identifying and ranking relevant item candidates. For instance, Gemini attains the highest P@1 values, indicating superior accuracy in recommending the top item. Additionally, Gemini’s strong performance in N@3 reflects its ability to rank relevant candidates effectively. GPT-3.5 and Mistral 8B models also demonstrate improvements over MPG. Both models show competitive results in P@3 and R@3, indicating their effectiveness in providing multiple relevant recommendations. GPT-3.5, for example, often comes close to Gemini’s performance, particularly in N@3, suggesting its capability to enhance the ranking quality significantly compared to MPG. The Mistral 8B model, while not always leading, maintains consistent and reliable performance across all metrics. Llama 3-8B, though slightly behind the other LLMs, still shows improvement over MPG in several metrics. While it may not be as effective in capturing the top preferences of users as the leading models, it provides a consistent alternative that surpasses the baseline.
(a) Keyword Order
(b) Candidate Order
(c) Position of Candidate
(d) Popularity Bias
Figure 7. Performance examination of LLMs across various aspects, conducted using Gemini Pro
RQ2. Is employing keywords more effective than using reviews in prompting? To convey user preferences and item characteristics, the straightforward approach is to use the full content of user reviews. However, due to the context length limitations in LLM prompts, KALM4REC uses keywords instead. We compare the effectiveness of LLMs when using reviews versus keywords for the re-ranking task on Gemini Pro. For each item candidate in a prompt, we describe it using a representative review, which is selected based on its distance to the review centroid. The centroid embedding is calculated as the average of the embeddings of all the item’s reviews. The comparison results across 3 cities are illustrated in Figure 5 in Precision@k and Recall@k with k = 1, 3. From Figure 5, we find that rankings by our KALM4REC consistently show remarkable improvements across all cities in P@1 and R@1. However, we observe KALM4REC with review gets a lower number in P@3 and R@3 compared to the retrieval model for Singapore. Whereas, using keywords offers improvements across all the metrics. This improvement is likely attributed to the detailed information provided by keywords, which directly correspond to the input keywords. This enables the LLM to better understand the relevance between the input user and the candidate items.
Another advantage of using keywords is their efficiency in prompting LLMs. Figure 6(a) indicates that prompting with keywords significantly reduces the total number of tokens compared to using reviews on Gemini Pro. This reduction in token length leads to lower associated costs and enhanced efficiency. Consequently, across all datasets, employing keywords not only offer superior effectiveness but also provide a more efficient and cost-effective approach for prompting LLMs.
RQ3. How do prompt design strategies impact the efficiency of LLMs? We compare the performance of KALM4REC with zero-shot [$\mathcal{H}_{z}$] and few-shot [$\mathcal{H}_{f}$] strategies. In this experiment, we set number of keywords for user ranging from 3 to 5 while number of keywords for item range from 3 to 20. We report the best setting with Gemini Pro for each strategy in Figure 6(b). We observe significant performance improvements across all strategies of zero-shot and few-shot when compared to MPG. The number of shots (1 shot, 2 shots, 3 shots) indicates the number of examples introduced to LLMs. Overall, increasing the number of examples enhances LLM performance, with the 3-shot approach mostly achieve the best results. This outcome is reasonable as LLMs can better capture user intent with more examples.
Figure 8. Comparison of the performance of various numbers of keywords used to describe candidates and users at Precision@1. The experiment is conducted using Gemini Pro on Yelp-Singapore data.
RQ4. Do LLMs suffer from biases while ranking? We adopt the insights from (Hou et al., 2024) that suggest several factors can affect the ranking performance of LLMs. To explore potential biases in LLM ranking, specifically on Gemini Pro, we conducted experiments focusing on various aspects. First, we examined how the order of keywords in the prompt affects performance. We shuffled the keywords for each testing user and reported the average performance of re-ranking. As shown in Figure 7a, the performance of re-ranking with shuffled keywords varies, indicating that keyword order influences LLMs. Notably, performance declines when keywords are not sorted by importance, suggesting that LLMs better understand user preferences when keywords are arranged according to their relevance. This effect is especially pronounced in Edinburgh, where performance with unordered keywords is even worse than with MPG.
Additionally, we investigated the impact of candidate order on LLMs. The results, shown in Figure 7b, reveal that Gemini Pro is sensitive to the order of candidates in the prompt. Shuffling the candidate order leads to a significant drop in baseline performance at P@3. However, LLMs’ re-ranking performance shows notable improvement. This suggests that even if the list of candidates is shuffled, LLMs can still recommend suitable items if ground truth items are presented. While the re-ranking performance with shuffled candidates is relatively stable, it remains lower compared to re-ranked MPG. This indicates that integrating efficient retrieval models with LLMs enhances overall performance, as the order of candidates by retrieval positively influences the results.
Next, we assessed LLM performance under the assumption that LLMs can recommend suitable items even if ground truth items are positioned in less favorable locations within the candidate list. We replaced all ground truth items in the candidate list at positions 5, 10, 15, and at the end, ensuring that each list contained no more than four ground truth items to avoid overlap. Results shown in Figure 7c indicate that LLMs are effective at identifying ground truth items even in challenging scenarios, demonstrating relatively stable performance across different settings. However, with the order provided by the retrieval module, KALM4REC achieved the best results highlighting the efficiency of our framework.
Finally, we explore how item popularity affects LLM recommendations. While item popularity cannot be measured directly, we use the normalized interaction count from the training set to represent it. Figure 7d illustrates the relationship between the average popularity score of items and their position in the re-ranked list (e.g., positions 0, 4, 9, 14). We find that highly popular items are often recommended at the top of the re-ranked list, whereas items with lower popularity tend to appear in lower positions. This suggests that popular items have a higher chance of being ranked favorably.
RQ5. The impact of the number of keywords on the re-ranking performance. Figure 8 presents heatmaps depicting the performance of KALM4REC across varying numbers of shots under different keyword conditions for users and items. The x-axis represents the number of keywords used to describe the user, while the y-axis represents the number of keywords used to describe the items. The color intensity reflects the model’s performance. The heatmaps reveal that increasing the number of shots generally enhances accuracy, fewer keywords can still yield relatively good performance. Besides, performance variations are observed depending on the number of keywords for users and items. We also acknowledge that the quality of keywords can affect KALM4REC’s performance. To investigate this, we compared various keyword extractors in Table 5. Our experiments show that SpaCy consistently delivered the best performance in keyword extraction.
Table 5. Comparison performance of our retrieval module with different keyword extractors
Keyword Extractors
spaCy Rake Yake NLTK
Dataset P@20 R@20 P@20 R@20 P@20 R@20 P@20 R@20
[…]
5. Discussion and Conclusion
In this work, we investigate the idea of augmenting LLMs for cold-start user recommendations with keywords extracted from user reviews. For retrieving potential candidates, we present CBR and MPG, keyword-based methods, which outperform conventional approaches. We then employ LLMs to re-rank the obtained candidates using designed prompting strategies that incorporate keywords to represent users and items. Comprehensive experiments indicate that KALM4REC is capable of handling cold-start user scenarios effectively. Additionally, our framework shows the potential of integrating different retrieval and language models to achieve promising performance in the future for multiple domains. Several factors may affect the performance of our framework. For instance, while the constructed graph in MPG is effective, the connections between keyword nodes should also be considered, as similar keywords may need to pass information to each other before reaching the restaurant nodes. However, implementing this could be computationally costly when the number of keywords is large. Additionally, the number of keywords introduced into the prompt may be reduced by increasing the number of example introductions, as shown in the experiments. For future work, we plan to develop technical approaches to enhance the robustness of each stage in our framework, leveraging the insights drawn from our comprehensive set of experiments.